### v1

# Exact fluctuation relation for open systems beyond the Jarzynski equality

Abstract
We derive exact fluctuation equalities for open systems that recover free energy differences between two equilibrium endpoints connected by nonequilibrium processes with arbitrary dynamics and coupling. The exponential of the free energy difference is expressed in terms of ensemble averages of the Hamiltonian of mean force (HMF) shift and the chi-squared divergence between the initial and final marginal probability distribution of the open system. A trajectory counterpart of this relation follows from an asymptotic equilibration postulate, which treats relaxation to the final stationary canonical state as a boundary condition rather than as a consequence of constraints on the driven dynamics. In the frozen-coupling regime, the HMF shift reduces to the bare-system Hamiltonian shift, yielding a clear heat–work decomposition. The Jarzynski equality (JE) is recovered under the assumption of Hamiltonian dynamics for the combined system. We validate the theory on a dissipative, phase-space-compressing drive followed by an underdamped Langevin relaxation, where the assumptions underlying the JE break down, whereas our equality reproduces the exact free energy differences.
preprint: APS/123-QED
The Jarzynski equality (JE) [1], $\langle e^{-\beta W}\rangle=e^{-\beta\Delta F}$, is a fundamental result in modern statistical mechanics. It establishes an exact connection between the exponential average of the nonequilibrium work $W$ performed during an irreversible process and the equilibrium free energy difference $\Delta F$. Along with related fluctuation theorems (FTs) [2, 3, 4], it has been confirmed in both classical and quantum systems and provides a powerful framework for extracting equilibrium information from nonequilibrium measurements [5, 6, 7, 8, 9, 10, 11, 12]. Nevertheless, its generality must be interpreted with care in different situations [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. The theoretical foundation of the original JE rests on strict dynamical reversibility [1]. In its original derivation, the composite system $\mathcal{S}+\mathcal{E}$, with bare system $\mathcal{S}$ and environment $\mathcal{E}$, evolves under deterministic dynamics that obey Liouville’s theorem, ensuring phase–space volume preservation. This preservation implies a fundamental dynamical constraint underlying the equality. A stochastic generalization soon followed, reformulating the dynamics through a Markov master equation or Langevin description [29]. In this generalization, reversibility takes a statistical form through detailed balance (DB). DB ensures that, for each fixed control parameter $\lambda$, the Gibbs–Boltzmann canonical distribution serves as the stationary state probability distribution, providing the probabilistic analogue of microscopic reversibility. The necessity of reversibility constraints was highlighted by Cohen and Mauzerall’s criticism [22], which questioned the formal validity of the JE and suggested its applicability might be limited to near-equilibrium weakly coupled regimes. Jarzynski’s reply [30] demonstrated that the JE is a mathematical identity following directly from Hamiltonian dynamics of the composite system $(\mathcal{S}+\mathcal{E})$. This derivation, later refined to incorporate strong coupling through the Hamiltonian of mean force (HMF) $H_{\beta}^{*}(X_{\mathcal{S}},\lambda)$, confirmed that while the free energy interpretation changes, the proof still relies on Liouville’s theorem and therefore on microscopic reversibility of the composite dynamics. Speck and Seifert [31] extended the equality to non-Markovian processes governed by the generalized Langevin equation. For such dynamics, the JE remains valid, provided the friction kernel and noise correlations obey the fluctuation–dissipation theorem (FDT). Compliance with FDT ensures that the Gibbs–Boltzmann distribution remains stationary for each frozen $\lambda$. While the JE demonstrated resilience against classical memory effects, a more fundamental challenge arose from an observation that the standard JE is necessarily violated in the presence of feedback — a phenomenon central to Maxwell’s Demon [32] and experimentally realized in information ratchets[33]. A formal resolution was achieved through the generalized Jarzynski equality (GJE) under feedback control, as formulated by Sagawa and Ueda [21, 14]. That extension incorporates the stochastic mutual information gained during measurement as a thermodynamic resource. However, this GJE still relies on a dynamical assumption, requiring the classical stochastic processes to satisfy local detailed balance (LDB) to ensure that the time-reversal property of path probabilities is well defined [34]. Regardless of the distinct approaches for the JE derivations and related generalizations, they all share an inherent dependency on underlying dynamical constraints such as microscopic reversibility, DB, FDT, or LDB [2, 35]. When these constraints are violated, the JE (or GJE) is no longer guaranteed to hold. Across biology, chemistry, physics, and engineering, many systems are driven under conditions that break these dynamical constraints, yet still relax to a genuine equilibrium once the driving is kept fixed. [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48].
The key question is whether it is possible to quantify the free energy difference between two equilibrium points connected by nonequilibrium processes, given arbitrary dynamics and system-environment coupling. In this Letter, we derive equalities for this most general scenario. As the first result, we derive an endpoint equality in terms of the driving and coupling protocols (Thm. 1), showing that its validity is independent of both the underlying dynamics and the strength of system–environment coupling linking the equilibrium endpoints. The second result (Thm. 2) establishes the trajectory counterpart based on the asymptotic postulate of thermal equilibration. For strongly coupled systems in the frozen–coupling regime — the physically most natural case — the equalities depend on the HMF in exactly the same way as on the bare–system Hamiltonian (Cor. 1). We then derive a heat–work decomposition (Cor. 2), which reveals the thermodynamic structure of the relations and provides a consistent interpretation of work and heat exchange. In the limit of Liouvillian dynamics and frozen coupling, our framework reduces to the standard JE. We validate the theory on a composite model, where the standard work estimator of JE fails under non-Liouvillian driving, while the endpoint and trajectory counterpart equalities reproduce the exact equilibrium free energy differences. We begin with a few necessary definitions (Def. 1-3) that later serve as the foundation for the endpoint and trajectory relations (Thm. 1 & 2) derived in this work.
Figure 1: (A) Schematic of the composite $\mathcal{S}+\mathcal{E}$ at a fixed bath temperature $T=\beta^{-1}/k_{B}$. The control $\lambda(t)$ (schematic purple spring) and the coupling $C(t)$ (schematic black spring) are driven independently and become constant at times $t_{\lambda}$ and $t_{C}$, respectively; the composite system then relaxes to equilibrium at $t_{\rm eq}$. The system free energy difference $\Delta F_{\mathcal{S}}^{*}(\beta)$ refers to the canonical endpoints $(\lambda(0),C(0))\!\to\!(\lambda(t_{\rm eq}),C(t_{\rm eq}))$ and is defined through the HMF partition functions. (B) An initial microstate $X_{0}\!\sim\!P_{0}(X,\beta)$ on $\Gamma(0)$ evolves under the kinematic map $X(t|X_{0})=\mathcal{T}_{t}(X_{0})$ through the ramp and the post–ramp relaxation, yielding the asymptotic density $P_{t_{\rm eq}}(X,\beta)$ on $\Gamma(t_{\rm eq})$. The frozen coupling case, $C(t)\equiv C(0)$, is the regime used in our validation.
Definition 1 (Microscopic trajectory of the composite system).
Here, we formulate the evolution of the composite system in an abstract manner, remaining agnostic to the details of its underlying dynamics. We consider a composite system $\mathcal{S}+\mathcal{E}$, consisting of a system of interest $\mathcal{S}$ and its environment $\mathcal{E}$. The total phase space of the composite at time $t$ is denoted by $\Gamma_{t}$, and a single microscopic state by $X\in\Gamma_{t}$. Each microstate contains the full phase-space coordinates of both parts, $X=(X_{\mathcal{S}},X_{\mathcal{E}})$, where $X_{\mathcal{S}}$ and $X_{\mathcal{E}}$ collect the positions and momenta of particles in $\mathcal{S}$ and $\mathcal{E}$, respectively. We introduce the trajectory map $\mathcal{T}_{t}:\Gamma_{0}\rightarrow\Gamma_{t}$, which assigns to each initial state $X_{0}\in\Gamma_{0}$, its evolved image $X(t|X_{0})=\mathcal{T}_{t}(X_{0})$ at time $t$. The map $\mathcal{T}_{t}$ is a purely kinematic construct and does not rely on any specific dynamical generator. It may represent Hamiltonian, stochastic, or arbitrary evolution, providing a unified description for all possible dynamics. Since $\mathcal{T}_{t}$ acts on the composite microstates $X_{0}\in\Gamma_{0}$, no assumption is made that the system and environment trajectories can be defined independently — a requirement intrinsic to any open system description. The microscopic state of the composite system along a trajectory is therefore written as $X(t|X_{0})=\bigl(X_{\mathcal{S}}(t|X_{0})=\mathcal{T}_{t}^{\mathcal{S}}(X_{0}),\,X_{\mathcal{E}}(t|X_{0})=\mathcal{T}_{t}^{\mathcal{E}}(X_{0})\bigr)$, where $\mathcal{T}_{t}^{\mathcal{S}}$ and $\mathcal{T}_{t}^{\mathcal{E}}$ denote the respective projections of the composite map $\mathcal{T}_{t}$ onto the system and environment coordinates. Next, we specify the Hamiltonian structure along the established trajectory.
Definition 2 (Total Hamiltonian along a trajectory).
The instantaneous total Hamiltonian of the composite system $\mathcal{S}+\mathcal{E}$ evaluated along the trajectory $\mathcal{T}_{t}(X_{0})$ is defined as
$\mathcal{H}_{\mathcal{S}+\mathcal{E}}\bigl(\mathcal{T}_{t}(X_{0}),\lambda(t),C(t)\bigr)=\mathcal{H}_{\mathcal{S}}\bigl(\mathcal{T}_{t}^{\mathcal{S}}(X_{0}),\lambda(t)\bigr)$
$\quad+\mathcal{H}_{\mathcal{E}}\bigl(\mathcal{T}_{t}^{\mathcal{E}}(X_{0})\bigr)+\mathcal{V}_{\mathcal{S}\mathcal{E}}\bigl(\mathcal{T}_{t}^{\mathcal{S}}(X_{0}),\mathcal{T}_{t}^{\mathcal{E}}(X_{0}),C(t)\bigr).$ (1)
Here $\mathcal{H}_{\mathcal{S}}$ and $\mathcal{H}_{\mathcal{E}}$ are the bare Hamiltonians of the system and environment, respectively, and $\mathcal{V}_{\mathcal{S}\mathcal{E}}$ is their interaction potential. Two externally controlled parameters appear: $\lambda(t)$ represents the driving protocol acting on $\mathcal{S}$ (for example, a control parameter or mechanical coordinate) and $C(t)$ controls the coupling strength or interaction channel between $\mathcal{S}$ and $\mathcal{E}$. By allowing $\lambda(t)$ and $C(t)$ to vary independently, we retain the most general energetic structure of an open system subject to simultaneously applied driving and coupling controls. Each protocol may correspond to a sudden quench or to a continuous ramp in time. We denote by $t_{\lambda}$ and $t_{c}$ the times at which the corresponding protocols become constant, $\lambda(t)=\mathrm{const}\quad\text{for }t\geq t_{\lambda}$ and $C(t)=\mathrm{const}\quad\text{for }t\geq t_{c}$. The composite system subsequently relaxes to equilibrium at a later time $t_{\mathrm{eq}}$. In general, $0\leq t_{\lambda},\;t_{c}\leq t_{\mathrm{eq}}$ such that $t_{\lambda}\neq t_{c}$, so the completion of driving, coupling, and equilibration do not necessarily coincide. This distinction will later be crucial when we analyze the frozen–coupling limit $C(t)\equiv C(0)$. Now, we introduce the canonical ensembles for the composite system and the system of interest $\mathcal{S}$. These equilibrium measures serve as statistical reference points for the initial and final endpoints of the process.
Definition 3 (Canonical ensembles of the composite system and marginal distributions of the system).
The equilibrium probability distribution of the composite system $\mathcal{S}+\mathcal{E}$ at temperature $\beta^{-1}/k_{\mathrm{B}}$ is defined in canonical form as $P(X,\beta)={e^{-\beta\,\mathcal{H}_{\mathcal{S}+\mathcal{E}}(X,\lambda,C)}}/{\mathcal{Z}_{\mathcal{S}+\mathcal{E}}(\lambda,C,\beta)},$ where $\mathcal{H}_{\mathcal{S}+\mathcal{E}}$ is the total Hamiltonian given in Eq. (2) and $\mathcal{Z}_{\mathcal{S}+\mathcal{E}}$ denotes the partition function of the composite system, $\mathcal{Z}_{\mathcal{S}+\mathcal{E}}(\lambda,C,\beta)=\int_{\Gamma_{t}}\mathrm{d}X\,e^{-\beta\,\mathcal{H}_{\mathcal{S}+\mathcal{E}}(X,\lambda,C)}.$ Two specific equilibrium ensembles are relevant to the present discussion, i.e.
$P_{0}(X,\beta)$ $=\frac{e^{-\beta\,\mathcal{H}_{\mathcal{S}+\mathcal{E}}(X,\lambda(0),C(0))}}{\mathcal{Z}_{\mathcal{S}+\mathcal{E}}(\lambda(0),C(0),\beta)},$ (2)
$P_{t_{\mathrm{eq}}}(X,\beta)$ $=\frac{e^{-\beta\,\mathcal{H}_{\mathcal{S}+\mathcal{E}}(X,\lambda(t_{\mathrm{eq}}),C(t_{\mathrm{eq}}))}}{\mathcal{Z}_{\mathcal{S}+\mathcal{E}}(\lambda(t_{\mathrm{eq}}),C(t_{\mathrm{eq}}),\beta)}.$ (3)
The distributions $P_{0}$ and $P_{t_{\mathrm{eq}}}$ describe the initial and final equilibrium states of the composite system at the same bath temperature $\beta^{-1}/k_{\mathrm{B}}$. The equilibrium marginal distribution of the system $\mathcal{S}$ is obtained by integrating over the environmental coordinates $P^{\mathcal{S}}(X_{\mathcal{S}},\beta)=\int\mathrm{d}X_{\mathcal{E}}\,P(X_{\mathcal{S}},X_{\mathcal{E}},\beta)$. This probability can be expressed in canonical form through the HMF, defined as
$\mathcal{H}_{\beta}^{*}(X_{\mathcal{S}},\lambda,C)=\mathcal{H}_{\mathcal{S}}(X_{\mathcal{S}},\lambda)$
$\qquad-\frac{1}{\beta}\ln\int\mathrm{d}X_{\mathcal{E}}\,e^{-\beta\,[\mathcal{H}_{\mathcal{E}}(X_{\mathcal{E}})+\mathcal{V}_{\mathcal{S}\mathcal{E}}(X_{\mathcal{S}},X_{\mathcal{E}},C)]}.$ (4)
Equation (3) defines an effective Hamiltonian that incorporates the influence of the environment on $\mathcal{S}$ at fixed $(\lambda,C,\beta)$. The corresponding partition function of the system is $\mathcal{Z}_{\mathcal{S}}^{*}(\lambda,C,\beta)=\int\mathrm{d}X_{\mathcal{S}}\,e^{-\beta\,\mathcal{H}_{\beta}^{*}(X_{\mathcal{S}},\lambda,C)}$. Hence the equilibrium marginal distribution of the system reads $P^{\mathcal{S}}(X_{\mathcal{S}},\beta)={e^{-\beta\,\mathcal{H}_{\beta}^{*}(X_{\mathcal{S}},\lambda,C)}}/{\mathcal{Z}_{\mathcal{S}}^{*}(\lambda,C,\beta)}$. Applying them to the two equilibrium endpoints of the process yields
$P^{\mathcal{S}}_{0}(X_{\mathcal{S}},\beta)$ $=\frac{e^{-\beta\,\mathcal{H}_{\beta}^{*}(X_{\mathcal{S}},\lambda(0),C(0))}}{\mathcal{Z}_{\mathcal{S}}^{*}(\lambda(0),C(0),\beta)},$ (5)
$P^{\mathcal{S}}_{t_{\mathrm{eq}}}(X_{\mathcal{S}},\beta)$ $=\frac{e^{-\beta\,\mathcal{H}_{\beta}^{*}(X_{\mathcal{S}},\lambda(t_{\mathrm{eq}}),C(t_{\mathrm{eq}}))}}{\mathcal{Z}_{\mathcal{S}}^{*}(\lambda(t_{\mathrm{eq}}),C(t_{\mathrm{eq}}),\beta)}.$ (6)
Eqs. (5) and (6) define the initial and final equilibrium probability distributions of the system in terms of its HMF, corresponding to the protocol endpoints $(\lambda(0),C(0))$ and $(\lambda(t_{\mathrm{eq}}),C(t_{\mathrm{eq}}))$, respectively. In Thm. 1, we show the exact connection between the system free energy difference at two endpoints and the canonical distributions and related HMFs at those endpoints.
Theorem 1 (Endpoint equalities for free energy differences).
For two equilibrium endpoints of the composite system $\mathcal{S}+\mathcal{E}$ prepared at the same inverse temperature $\beta$, we have (see Appendix for the proof)
$e^{-\beta\Delta F^{*}_{\mathcal{S}}(\beta)}=\frac{\bigl\langle e^{-\beta\Delta\mathcal{H}^{*}_{\beta}(X_{\mathcal{S}})}\bigr\rangle_{\mathcal{S}}}{1+\chi^{2}\bigl(P^{\mathcal{S}}_{t_{\mathrm{eq}}}\parallel P^{\mathcal{S}}_{0}\bigr)},$ (7)
and
$e^{+\beta\Delta F^{*}_{\mathcal{S}}(\beta)}=\bigl\langle e^{+\beta\Delta\mathcal{H}^{*}_{\beta}(X_{\mathcal{S}})}\bigr\rangle_{\mathcal{S}}.$ (8)
The free energy difference is
$\Delta F^{*}_{\mathcal{S}}(\beta)=F^{*}_{\mathcal{S}}(\lambda(t_{\mathrm{eq}}),C(t_{\mathrm{eq}}),\beta)-F^{*}_{\mathcal{S}}(\lambda(0),C(0),\beta),$ (9)
where $F^{*}_{\mathcal{S}}(\lambda,C,\beta)=\beta^{-1}\ln\mathcal{Z}_{\mathcal{S}}^{*}(\lambda,C,\beta)$ and the HMF shift is
$\Delta\mathcal{H}^{*}_{\beta}(X_{\mathcal{S}})=\mathcal{H}^{*}_{\beta}(X_{\mathcal{S}},\lambda(t_{\mathrm{eq}}),C(t_{\mathrm{eq}}))$
$\qquad-\mathcal{H}^{*}_{\beta}(X_{\mathcal{S}},\lambda(0),C(0)).$ (10)
Averages are taken over the final equilibrium ensemble,
$\langle\bullet\rangle_{\mathcal{S}}=\int\mathrm{d}X_{\mathcal{S}}\,\bullet\,P^{\mathcal{S}}_{t_{\mathrm{eq}}}(X_{\mathcal{S}},\beta),$ (11)
and the chi–squared divergence between the endpoint marginals is
$1+\chi^{2}\bigl(P^{\mathcal{S}}_{t_{\mathrm{eq}}}\parallel P^{\mathcal{S}}_{0}\bigr)=\int\mathrm{d}X_{\mathcal{S}}\,\frac{\bigl(P^{\mathcal{S}}_{t_{\mathrm{eq}}}(X_{\mathcal{S}},\beta)\bigr)^{2}}{P^{\mathcal{S}}_{0}(X_{\mathcal{S}},\beta)}.$ (12)
The result established here coincides with the endpoint equality derived in  [49], where the protocol is implicitly present. As can be seen from the derivation of Eqs. (7) and (8), no restriction is imposed on either the coupling or the form of the dynamics. The reasoning relies only on the existence of well-defined equilibrium endpoints, which guarantees that the corresponding marginal distributions and free energy difference are statistically and thermodynamically meaningful, respectively. Before introducing the trajectory counterparts, it is essential to clarify the assumption underlying their derivation. We postulate asymptotic equilibration of the full probability distribution along the trajectory. This assumption expresses the natural requirement that the Helmholtz free energy difference $\Delta F$ be well defined from the trajectory perspective of any nonequilibrium process connecting two equilibrium states, without imposing any particular dynamical constraint during the driving stage. It contrasts sharply with the standard derivations of the JE and its extensions, where the equality holds only if the underlying dynamics, during the application of the driving protocol, evolve toward a canonical stationary state once the control parameters are held fixed. In other words, the stationary canonical distribution at equilibrium is a consequence of constraints on the driven dynamics. By treating asymptotic convergence to $P_{t_{\mathrm{eq}}}$ as a natural boundary condition of thermal contact — rather than as a consequence of these dynamical constraints — the definition of equilibrium becomes decoupled from the specific generator of dynamics during the driving stage. This decoupling extends the application of our derivations to a broader class of systems, including those governed by arbitrary dynamics that are incompatible with the underlying assumptions of JE. We will later discuss the subtleties surrounding the imposition of that boundary condition in the validation section.
Theorem 2 (Trajectory counterpart of the endpoint equalities).
Using the trajectory map $\mathcal{T}_{t}:\Gamma_{0}\rightarrow\Gamma_{t}$ defined in Def. 1, the time evolution of the probability density of the composite system can be defined as
$P_{t}(X,\beta)=\int_{\Gamma_{0}}\mathrm{d}X_{0}\,\delta\bigl(X-\mathcal{T}_{t}(X_{0})\bigr)\,P_{0}(X_{0},\beta),$ (13)
where $P_{0}(X_{0},\beta)$ is the initial canonical distribution given in Eq. (2). Integrating over the environmental degrees of freedom yields the time evolution of the system’s marginal distribution (see Appendix for derivation)
$P_{t}^{\mathcal{S}}(X_{\mathcal{S}},\beta)=\int_{\Gamma_{0}}\mathrm{d}X_{0}\,P_{0}(X_{0},\beta)\,\delta\bigl(X_{\mathcal{S}}-\mathcal{T}_{t}^{\mathcal{S}}(X_{0})\bigr).$ (14)
We postulate the asymptotic equilibration of the full distribution that implies convergence of the time evolution of the system’s marginal distribution to its canonical endpoint form
$\lim_{t\rightarrow t_{\mathrm{eq}}}P_{t}^{\mathcal{S}}(X_{\mathcal{S}},\beta)=P^{\mathcal{S}}_{t_{\mathrm{eq}}}(X_{\mathcal{S}},\beta).$ (15)
Under that condition, the ensemble averages in Thm. 1 admit the exact trajectory representation (see Appendix for proof), with averages over the initial canonical ensemble, $\langle\bullet\rangle_{X_{0}}=\int\mathrm{d}X_{0}\,\bullet\,P_{0}(X_{0},\beta)$. Consequently, the endpoint equalities (7) and (8) take the trajectory form
$e^{-\beta\Delta F^{*}_{\mathcal{S}}(\beta)}=\frac{\bigl\langle e^{-\beta\Delta\mathcal{H}^{*}_{\beta}\bigl(\mathcal{T}_{t_{\mathrm{eq}}}^{\mathcal{S}}(X_{0})\bigr)}\bigr\rangle_{X_{0}}}{1+\chi^{2}\bigl(P^{\mathcal{S}}_{t_{\mathrm{eq}}}\parallel P^{\mathcal{S}}_{0}\bigr)},$ (16)
and
$e^{+\beta\Delta F^{*}_{\mathcal{S}}(\beta)}=\Bigl\langle e^{+\beta\Delta\mathcal{H}^{*}_{\beta}\bigl(\mathcal{T}_{t_{\mathrm{eq}}}^{\mathcal{S}}(X_{0})\bigr)}\Bigr\rangle_{X_{0}}.$ (17)
Frozen coupling regime. The scenario of frozen coupling deserves particular attention. In most free energy difference calculations — both theoretical and experimental — the system remains in fixed coupling (the coupling strength $C$ between $\mathcal{S}$ and $\mathcal{E}$ is kept constant throughout the process) with its environment while the control parameter $\lambda(t)$ is driven. This is precisely the regime underlying the standard derivation of the JE, although the JE itself is agnostic to whether the protocol acts as a driving or a coupling. Within this common, physically motivated freezing scenario, the formal structure of our derivation simplifies considerably.
Corollary 1 (Frozen coupling: trajectory forms reduce to bare-system expressions).
For a fixed coupling protocol $C(t)\equiv C(0)$, the HMF increment along a trajectory reduces exactly to the difference of bare–system Hamiltonians,
$\Delta\mathcal{H}^{*}_{\beta}\bigl(\mathcal{T}_{t_{\rm eq}}^{\mathcal{S}}(X_{0})\bigr)$
$=\mathcal{H}_{\beta}^{*}\bigl(\mathcal{T}_{t_{\rm eq}}^{\mathcal{S}}(X_{0}),\lambda(t_{\rm eq}),C(0)\bigr)-\mathcal{H}_{\beta}^{*}\bigl(\mathcal{T}_{t_{\rm eq}}^{\mathcal{S}}(X_{0}),\lambda(0),C(0)\bigr)$
$=\mathcal{H}_{\mathcal{S}}\bigl(\mathcal{T}_{t_{\rm eq}}^{\mathcal{S}}(X_{0}),\lambda(t_{\rm eq})\bigr)-\mathcal{H}_{\mathcal{S}}\bigl(\mathcal{T}_{t_{\rm eq}}^{\mathcal{S}}(X_{0}),\lambda(0)\bigr),$ (18)
Equation (1) follows directly from Eq. (3), for constant $C$, the environmental contribution inside the logarithmic term cancels between the two endpoints. Substituting Eq. (1) into the trajectory equalities (16)–(17) yields
$e^{-\beta\Delta F^{*}_{\mathcal{S}}(\beta)}$
$=\frac{\Big\langle e^{-\beta\mathcal{H}_{\mathcal{S}}\bigl(\mathcal{T}_{t_{\mathrm{eq}}}^{\mathcal{S}}(X_{0}),\lambda(t_{\mathrm{eq}})\bigr)}e^{+\beta\mathcal{H}_{\mathcal{S}}\bigl(\mathcal{T}_{t_{\mathrm{eq}}}^{\mathcal{S}}(X_{0}),\lambda(0)\bigr)}\Big\rangle_{X_{0}}}{1+\chi^{2}\bigl(P^{\mathcal{S}}_{t_{\mathrm{eq}}}\parallel P^{\mathcal{S}}_{0}\bigr)},$ (19)
and the complementary identity,
$e^{+\beta\Delta F^{*}_{\mathcal{S}}(\beta)}$
$=\Big\langle e^{+\beta\mathcal{H}_{\mathcal{S}}\bigl(\mathcal{T}_{t_{\mathrm{eq}}}^{\mathcal{S}}(X_{0}),\lambda(t_{\mathrm{eq}})\bigr)}e^{-\beta\mathcal{H}_{\mathcal{S}}\bigl(\mathcal{T}_{t_{\mathrm{eq}}}^{\mathcal{S}}(X_{0}),\lambda(0)\bigr)}\Big\rangle_{X_{0}}.$ (20)
It is important to mention that as soon as $C(t)$ is non-frozen, the cancellation in (1) no longer occurs and the full HMF structure must be retained.
Corollary 2 (Heat–work decomposition for frozen coupling).
We now discuss the thermodynamic structure hidden in Eqs. (1) and  (1). In the frozen coupling regimes, energetic changes in $\mathcal{S}$ arise from (i) external manipulation of $\lambda(t)$ and (ii) exchange with the environment through the microscopic evolution of $X_{\mathcal{E}}$. To inspect these contributions, we start from Eq. (1) and apply a simple algebraic insertion–subtraction of the initial energy at fixed protocol, $\mathcal{H}_{\mathcal{S}}\bigl(X_{\mathcal{S}}(0|X_{0}),\lambda(0)\bigr)$, which yields the following representation. For $C(t)\equiv C(0)$,
$\mathcal{H}_{\beta}^{*}\bigl(\mathcal{T}_{t_{\rm eq}}^{\mathcal{S}}(X_{0}),\lambda(t_{\rm eq}),C(0)\bigr)-\mathcal{H}_{\beta}^{*}\bigl(\mathcal{T}_{t_{\rm eq}}^{\mathcal{S}}(X_{0}),\lambda(0),C(0)\bigr)$
$=\underbrace{\Big[\mathcal{H}_{\mathcal{S}}\bigl(X_{\mathcal{S}}(t_{\mathrm{eq}}|X_{0}),\lambda(t_{\mathrm{eq}})\bigr)-\mathcal{H}_{\mathcal{S}}\bigl(X_{\mathcal{S}}(0|X_{0}),\lambda(0)\bigr)\Big]}_{I(t_{\mathrm{eq}}|X_{0})}$
$-\underbrace{\Big[\mathcal{H}_{\mathcal{S}}\bigl(X_{\mathcal{S}}(t_{\mathrm{eq}}|X_{0}),\lambda(0)\bigr)-\mathcal{H}_{\mathcal{S}}\bigl(X_{\mathcal{S}}(0|X_{0}),\lambda(0)\bigr)\Big]}_{II(t_{\mathrm{eq}}|X_{0})}.$ (21)
Term $I$. Along each realization, the system follows the trajectory $X_{\mathcal{S}}(t|X_{0})$ in its phase space. The total time variation of the bare–system Hamiltonian $\mathcal{H}_{\mathcal{S}}(X_{\mathcal{S}}(t|X_{0}),\lambda(t))$ is obtained by applying the chain rule to its explicit dependence on both $X_{\mathcal{S}}$ and the control parameter $\lambda(t)$,
$\frac{\mathrm{d}}{\mathrm{d}t}\,\mathcal{H}_{\mathcal{S}}\bigl(X_{\mathcal{S}}(t|X_{0}),\lambda(t)\bigr)=\underbrace{\bigl(\nabla_{X_{\mathcal{S}}}\mathcal{H}_{\mathcal{S}}\bigr)\cdot\dot{X}_{\mathcal{S}}}_{\dot{Q}_{\mathcal{S}}}+\underbrace{\frac{\partial\mathcal{H}_{\mathcal{S}}}{\partial\lambda}\,\dot{\lambda}}_{\dot{W}_{\mathcal{S}}}.$ (22)
The first term describes the energy exchange with the environment along the actual phase–space trajectory and defines the instantaneous heat rate, $\dot{Q}_{\mathcal{S}}$, whereas the second term accounts for the parametric energy input due to the protocol and defines the instantaneous work rate, $\dot{W}_{\mathcal{S}}$. Following the standard conventions of stochastic energetics [50], their time integrals give
$W_{\mathcal{S}}(t_{\mathrm{eq}}|X_{0})$ $=\int_{0}^{t_{\mathrm{eq}}}{\rm d}t\,\frac{\partial\mathcal{H}_{\mathcal{S}}}{\partial\lambda}\,\dot{\lambda}(t),$ (23)
$Q_{\mathcal{S}}(t_{\mathrm{eq}}|X_{0})$ $=\int_{0}^{t_{\mathrm{eq}}}{\rm d}t\,\big(\nabla_{X_{\mathcal{S}}}\mathcal{H}_{\mathcal{S}}\big)\cdot\dot{X}_{\mathcal{S}}(t|X_{0}),$ (24)
so that integrating Eq. (22) over $t\in[0,t_{\mathrm{eq}}]$ yields the total increment
$I(t_{\mathrm{eq}}|X_{0})=W_{\mathcal{S}}(t_{\mathrm{eq}}|X_{0})+Q_{\mathcal{S}}(t_{\mathrm{eq}}|X_{0}).$ (25)
Term $II$. For a fixed control $\lambda(0)$, $II(t_{\mathrm{eq}}|X_{0})$ can be written as
$II(t_{\mathrm{eq}}|X_{0})=\int_{0}^{t_{\mathrm{eq}}}\mathrm{d}t\;\Big(\nabla_{X_{\mathcal{S}}}\mathcal{H}_{\mathcal{S}}(X_{\mathcal{S}},\lambda(0))\Big)\cdot\dot{X}_{\mathcal{S}}.$ (26)
Here the gradient is evaluated with respect to $\mathcal{H}_{\mathcal{S}}$ frozen at $\lambda(0)$, whereas the $\dot{X}_{\mathcal{S}}$ belongs to the driven trajectory generated by $\lambda(t)$. Consequently, $II$ is not a thermodynamic heat unless $\dot{\lambda}=0$ for all $t$; it is a reference functional that projects the driven generalized velocity ($\dot{X}_{\mathcal{S}}$) onto the generalized force field ($\nabla_{X_{\mathcal{S}}}\mathcal{H}_{\mathcal{S}}$) of the initial stationary protocol. Because the reference and realized dynamics belong to distinct protocol layers, $II$ acts as a feedback-like correction without a definite thermodynamic notion. Combining Eqs. (25) and (26) with Eq. (2), the total bare–system increment becomes
$W_{\mathcal{S}}(t_{\mathrm{eq}}|X_{0})+Q_{\mathcal{S}}(t_{\mathrm{eq}}|X_{0})-II(t_{\mathrm{eq}}|X_{0}).$ (27)
Substituting Eq. (27) into Eqs. (1)–(1) yields the compact heat–work representation of the endpoint equalities,
$e^{-\beta\Delta F_{\mathcal{S}}^{*}(\beta)}$ $=\frac{\big\langle e^{-\beta[\,W_{\mathcal{S}}+Q_{\mathcal{S}}-II\,]}\big\rangle_{X_{0}}}{1+\chi^{2}(P^{\mathcal{S}}_{t_{\mathrm{eq}}}\parallel P^{\mathcal{S}}_{0})},$ (28)
$e^{+\beta\Delta F_{\mathcal{S}}^{*}(\beta)}$ $=\big\langle e^{+\beta[\,W_{\mathcal{S}}+Q_{\mathcal{S}}-II\,]}\big\rangle_{X_{0}}.$ (29)
The decomposition isolates three pathwise contributions: (i) $W_{\mathcal{S}}$, mechanical work due to driving; (ii) $Q_{\mathcal{S}}$, total heat exchanged during the whole process; and (iii) $II$, a reference (projection) functional obtained by contracting the driven velocity ($\dot{X}_{\mathcal{S}}$) with the $\lambda(0)$ force field ($\nabla_{X_{\mathcal{S}}}\mathcal{H}_{\mathcal{S}}$). Only when $\lambda$ is held fixed does $II$ coincide with a thermodynamic heat.
GJE connection. The GJE restores $e^{-\beta\Delta F}$ under feedback by augmenting work with mutual information, providing an operational measure of information as a thermodynamic resource [21]. Our identities achieve restoration without invoking measurement–feedback a priori and without assuming LDB. The pathwise correction $Q_{\mathcal{S}}-II$ compensates dynamical asymmetry during driving, and the ensemble overlap $1+\chi^{2}(P^{\mathcal{S}}_{t_{\rm eq}}\parallel P^{\mathcal{S}}_{0})$ compensates endpoint mismatch. Formally, our equality is broader in scope, it holds without feedback and without LDB, while its structure is fully compatible with the GJE.
Figure 2: (A) System free energy difference $\Delta F_{\mathcal{S}}^{*}$ versus the final control value $\lambda(t_{\rm eq})$: circles (JE), squares (Eq. 1), triangles (Eq. 1), solid line (exact). Ramp/post-ramp protocol: $\lambda(t)$ increases linearly from $\lambda(0)$ to the target over $t_{\lambda}$, then is held fixed. (B) Representative run for $\lambda(0)=1\!\to\!\lambda(t_{\rm eq})=4$. Center: $\lambda(t)$. Top/bottom: ensemble trajectories of $x(t)$ and $p_{x}(t)$ (dashed line marks $t_{\lambda}$). Left/right: canonical marginals $P^{\mathcal{S}}_{0}(x,p_{x})$ and $P^{\mathcal{S}}_{t_{\rm eq}}(x,p_{x})$. During the ramp, dynamics are non-Liouvillian with damping $(\gamma_{\rm nl}^{\mathcal{S}},\gamma_{\rm nl}^{\mathcal{E}})$, post-ramp relaxation is governed by the underdamped Langevin equations, whose solution converges to the canonical distribution of the composite system, thereby satisfying the asymptotic equilibration postulate, Eq. (15).
JE as limiting case. The central identity underlying all our derivations is
$e^{-\beta\Delta F_{\mathcal{S}}^{*}}=\frac{Z^{*}(\lambda(t_{\mathrm{eq}}),C(t_{\mathrm{eq}}),\beta)}{Z^{*}(\lambda(0),C(0),\beta)}.$ (30)
By the definition of the HMF, $Z_{\mathcal{S}+\mathcal{E}}(\lambda,C,\beta)=Z^{*}(\lambda,C,\beta)\,Z_{\mathcal{E}}(\beta)$, so Eq. (30) is equivalent to
$\frac{Z^{*}(\lambda(t_{\mathrm{eq}}),C(t_{\mathrm{eq}}),\beta)}{Z^{*}(\lambda(0),C(0),\beta)}=\frac{Z_{\mathcal{S}+\mathcal{E}}(\lambda(t_{\mathrm{eq}}),C(t_{\mathrm{eq}}),\beta)}{Z_{\mathcal{S}+\mathcal{E}}(\lambda(0),C(0),\beta)}.$ (31)
This equality, which corresponds to [Eq. (23) in [30]], contains no reference to the dynamics of the composite system. All relations derived in this work—independent of coupling strength or dynamical assumptions—reduce to this equilibrium partition function ratio. The standard JE can be viewed as the specific realization of this identity under Hamiltonian evolution of the composite system. For completeness, the detailed derivation linking Eq. (31) to the conventional JE is provided in Appendix, where we recast Jarzynski’s original argument using the present notation.
Validation. To validate the framework, we examine how the proposed equalities hold beyond the regime where the JE applies. The model construction is inspired by the two–stage thought experiment introduced by Jarzynski in his reply [30] to Cohen’s critique [22]. In his argument, the control parameter is first driven from $\lambda(0)$ to $\lambda(t_{\lambda})$ (stage 1) and then held fixed at $\lambda=\lambda(t_{\lambda})$ for relaxation (stage 2). Two observers — one stopping at $t_{\lambda}$ (out of equilibrium) and the other at $t_{\mathrm{eq}}$ (after relaxation) — record the same work, since no work is performed once the driving protocol is fixed, and calculate the same $\Delta F_{\mathcal{S}}^{*}$, defined as the equilibrium free energy difference between the canonical states associated with $\lambda(0)$ and $\lambda(t_{\lambda})=\lambda(t_{\mathrm{eq}})$ via the HMF partition functions. As Jarzynski emphasized [30], ”whether or not we choose to include a relaxation stage has no bearing on the validity” of the JE. While this statement is entirely correct, it implicitly depends on a key assumption about the underlying dynamics. To probe this assumption directly, we construct a validation model consisting of a non–Liouvillian ramp (stage 1) that breaks phase–space volume preservation, followed by an underdamped Langevin relaxation (stage 2) that enforces the required equilibrium endpoint. In this scenario, the JE no longer holds because phase–space contraction during the ramp breaks the Liouvillian structure of the composite dynamics, preventing the equilibrium partition function ratio in Eq. (31) from reproducing the JE. For a detailed discussion of the validation model and protocol, see the Appendix.
Conclusions: In this Letter, we establish exact fluctuation relations for open systems. Using the HMF, the free energy difference is expressed via exponential moments of the HMF shift and an endpoint chi-squared divergence (Thm. 1). The trajectory counterparts of that relation is obtained under the assumption of asymptotic equilibration (Thm. 2). In the frozen coupling regime the relations reduce to bare-system expressions (Cor. 1) with a clear heat–work decomposition (Cor. 2), and they include the JE and its feedback extensions (GJE) as limiting cases. Validation on a composite model with a non-Liouvillian ramp followed by underdamped Langevin relaxation shows that our endpoint and trajectory relations reproduce the exact free energy differences, demonstrating accuracy and practical applicability beyond the JE regime. The generality of this framework invites direct application to complex nonequilibrium settings where traditional FTs are fragile or inapplicable, including active matter [51, 52], biological processes [53], complex molecular and chemical environments [54], and open quantum systems [55]. In these regimes, our equalities provide a practical route to reconstruct free energy landscapes from realistic driving protocols without imposing idealized dynamical models or weak coupling assumptions. When applied to situations in which the JE is valid, our relations reduce to an exact re-expression of the JE that makes explicit the roles of endpoint overlap and system–environment interaction through the chi-squared factor and the HMF structure. In this sense, the JE becomes a diagnostic tool: deviations can be traced to insufficient overlap [56], rare-event sampling issues [57], or neglected coupling effects [50]. These diagnostics offer concrete guidance for designing and refining experimental and simulation protocols [58], and we expect them to support more reliable free energy estimation across a broad class of strongly driven open systems.
Acknowledgements.
We gratefully acknowledge financial support by the DFG under Germany’s Excellence Strategy EXC 2089/1-390776260 (e-conversion).