### v2

# Exact fluctuation relation for open systems beyond the Jarzynski equality

Abstract
We derive exact fluctuation equalities for open systems that recover free energy differences between two equilibrium endpoints connected by nonequilibrium processes with arbitrary dynamics and coupling. The exponential of the free energy difference is expressed in terms of ensemble averages of the Hamiltonian of mean force (HMF) shift and the chi-squared divergence between the initial and final marginal probability distribution of the open system. A trajectory counterpart of this relation follows from an asymptotic equilibration postulate, which treats relaxation to the final stationary canonical state as a boundary condition rather than as a consequence of constraints on the driven dynamics. In the frozen-coupling regime, the HMF shift reduces to the bare-system Hamiltonian shift, yielding a clear heat-work-feedback decomposition. The Jarzynski equality (JE) is recovered under the assumption of Hamiltonian dynamics for the combined system. We validate the theory on a dissipative, phase-space-compressing drive followed by an underdamped Langevin relaxation, where the assumptions underlying the JE break down, whereas our equality reproduces the exact free energy differences.
preprint: APS/123-QED
The Jarzynski equality (JE) [jarzynski1997nonequilibrium], $\langle e^{-\beta W}\rangle=e^{-\beta\Delta F}$, is a fundamental result in modern statistical mechanics. It establishes an exact connection between the exponential average of the nonequilibrium work $W$ performed during an irreversible process and the equilibrium free energy difference $\Delta F$. Along with related fluctuation theorems (FTs) [esposito2009nonequilibrium, crooks1998nonequilibrium, crooks1999entropy], it has been confirmed in both classical and quantum systems and provides a powerful framework for extracting equilibrium information from nonequilibrium measurements [liphardt2002equilibrium, ohzeki2010quantum, xiong2018experimental, liu2023experimental, harris2007experimental, douarche2005experimental, saira2012test, wimsatt2021harnessing]. Nevertheless, its generality must be interpreted with care in different situations [gore2003bias, toyabe2010experimental, deng2017deformed, hernandez2023experimental, hahn2023quantum, presse2006ordering, chen2008crooks, crooks2009comment, sagawa2010generalized, cohen2004note, cohen2005jarzynski, argun2016experimental, vilar2008failure, palmieri2007jarzynski, monge2018experimental, sone2020quantum]. The theoretical foundation of the original JE rests on strict dynamical reversibility [jarzynski1997nonequilibrium]. In its original derivation, the composite system $\mathcal{S}+\mathcal{E}$, with bare system $\mathcal{S}$ and environment $\mathcal{E}$, evolves under deterministic dynamics that obey Liouville’s theorem, ensuring phase-space volume preservation. This preservation implies a fundamental dynamical constraint underlying the equality. A stochastic generalization soon followed, reformulating the dynamics through a Markov master equation or Langevin description [jarzynski1997equilibrium]. In this generalization, reversibility takes a statistical form through detailed balance (DB). DB ensures that, for each fixed control parameter $\lambda$, the Gibbs-Boltzmann canonical distribution serves as the stationary state probability distribution, providing the probabilistic analogue of microscopic reversibility. The necessity of reversibility constraints was highlighted by Cohen and Mauzerall’s criticism [cohen2004note], which questioned the formal validity of the JE and suggested its applicability might be limited to near-equilibrium weakly coupled regimes. Jarzynski’s reply [jarzynski2004nonequilibrium] demonstrated that the JE is a mathematical identity following directly from Hamiltonian dynamics of the composite system $(\mathcal{S}+\mathcal{E})$. This derivation, later refined to incorporate strong coupling through the Hamiltonian of mean force (HMF) $H_{\beta}^{*}(X_{\mathcal{S}},\lambda)$, confirmed that while the free energy interpretation changes, the proof still relies on Liouville’s theorem and therefore on microscopic reversibility of the composite dynamics. Speck and Seifert [speck2007jarzynski] extended the equality to non-Markovian processes governed by the generalized Langevin equation. For such dynamics, the JE remains valid, provided the friction kernel and noise correlations obey the fluctuation-dissipation theorem (FDT). Compliance with FDT ensures that the Gibbs-Boltzmann distribution remains stationary for each frozen $\lambda$. While the JE demonstrated resilience against classical memory effects, a more fundamental challenge arose from an observation that the standard JE is necessarily violated in the presence of feedback — a phenomenon central to Maxwell’s Demon [maruyama2009colloquium] and experimentally realized in information ratchets [serreli2007molecular]. A formal resolution was achieved through the generalized Jarzynski equality (GJE) under feedback control, as formulated by Sagawa and Ueda [sagawa2010generalized, toyabe2010experimental]. That extension incorporates the stochastic mutual information gained during measurement as a thermodynamic resource. However, this GJE still relies on a dynamical assumption, requiring the classical stochastic processes to satisfy local detailed balance (LDB) to ensure that the time-reversal property of path probabilities is well defined [maes2021local]. Regardless of the distinct approaches for the JE derivations and related generalizations, they all share an inherent dependency on underlying dynamical constraints such as microscopic reversibility, DB, FDT, or LDB [esposito2009nonequilibrium, scholl2006proof]. When these constraints are violated, the JE (or GJE) is no longer guaranteed to hold. Across biology, chemistry, physics, and engineering, many systems are driven under conditions that break these dynamical constraints, yet still relax to a genuine equilibrium once the driving is kept fixed. [gaspard2020active, martinez2019inferring, gnesotto2018broken, goswami2019work, godec2022challenges, krishnamurthy2016micrometre, reimann2002brownian, bechinger2016active, caprini2019activity, vicsek2012collective, cates2015motility, fodor2016far, argun2016non]. The key question is whether it is possible to quantify the free energy difference between two equilibrium points connected by nonequilibrium processes, given arbitrary dynamics and system-environment coupling. A detailed treatment is given in a companion article [PRE], here we only recall the main setting needed for the present Letter. We consider a composite system $\mathcal{S}+\mathcal{E}$ with microscopic states $X=(X_{\mathcal{S}},X_{\mathcal{E}})\in\Gamma_{t}$. Its evolution is described by a trajectory map $\mathcal{T}_{t}:\Gamma_{0}\to\Gamma_{t}$,
$X(t|X_{0})=\mathcal{T}_{t}(X_{0}),$ (1)
with projections $\mathcal{T}_{t}^{\mathcal{S}}$, $\mathcal{T}_{t}^{\mathcal{E}}$ onto system and environment coordinates. The map $\mathcal{T}_{t}$ is purely kinematic and may represent Hamiltonian, stochastic, or more general dynamics; no specific generator is imposed (see [PRE] for details). Along a trajectory the total Hamiltonian is
$\mathcal{H}_{\mathcal{S}+\mathcal{E}}\bigl(\mathcal{T}_{t}(X_{0}),\lambda(t),C(t)\bigr)=\mathcal{H}_{\mathcal{S}}\bigl(\mathcal{T}_{t}^{\mathcal{S}}(X_{0}),\lambda(t)\bigr)$
$\quad+\mathcal{H}_{\mathcal{E}}\bigl(\mathcal{T}_{t}^{\mathcal{E}}(X_{0})\bigr)+\mathcal{V}_{\mathcal{S}\mathcal{E}}\bigl(\mathcal{T}_{t}^{\mathcal{S}}(X_{0}),\mathcal{T}_{t}^{\mathcal{E}}(X_{0}),C(t)\bigr),$ (2)
where $\mathcal{H}_{\mathcal{S}}$, $\mathcal{H}_{\mathcal{E}}$, and $\mathcal{V}_{\mathcal{S}\mathcal{E}}$ are the bare system, environment, and interaction contributions. The protocol $\lambda(t)$ drives $\mathcal{S}$, while $C(t)$ controls its coupling to $\mathcal{E}$. We denote by $t_{\lambda}$ and $t_{c}$ the times after which $\lambda(t)$ and $C(t)$ are fixed, and by $t_{\mathrm{eq}}$ the later time at which the composite has relaxed to equilibrium at $(\lambda(t_{\mathrm{eq}}),C(t_{\mathrm{eq}}))$. At initial and final fixed $(\lambda,C,\beta)$ the composite is in the canonical state
$P(X;\lambda,C,\beta)=\frac{e^{-\beta\,\mathcal{H}_{\mathcal{S}+\mathcal{E}}(X,\lambda,C)}}{\mathcal{Z}_{\mathcal{S}+\mathcal{E}}(\lambda,C,\beta)},$ (3)
with partition function
$\mathcal{Z}_{\mathcal{S}+\mathcal{E}}(\lambda,C,\beta)=\int_{\Gamma_{t}}\mathrm{d}X\,e^{-\beta\,\mathcal{H}_{\mathcal{S}+\mathcal{E}}(X,\lambda,C)}.$ (4)
The two equilibrium endpoints are
$P_{0}(X,\beta)$ $=P\bigl(X;\lambda(0),C(0),\beta\bigr),$ (5)
$P_{t_{\mathrm{eq}}}(X,\beta)$ $=P\bigl(X;\lambda(t_{\mathrm{eq}}),C(t_{\mathrm{eq}}),\beta\bigr).$ (6)
The equilibrium marginal of the system is
$P^{\mathcal{S}}(X_{\mathcal{S}};\lambda,C,\beta)=\int\mathrm{d}X_{\mathcal{E}}\,P(X_{\mathcal{S}},X_{\mathcal{E}};\lambda,C,\beta).$ (7)
It admits a canonical representation through the Hamiltonian of mean force (HMF)
$\mathcal{H}_{\beta}^{*}(X_{\mathcal{S}},\lambda,C)=\mathcal{H}_{\mathcal{S}}(X_{\mathcal{S}},\lambda)$
$-\frac{1}{\beta}\ln\int\mathrm{d}X_{\mathcal{E}}\,\frac{1}{\mathcal{Z}_{\mathcal{E}}}e^{-\beta\,[\mathcal{H}_{\mathcal{E}}(X_{\mathcal{E}})+\mathcal{V}_{\mathcal{S}\mathcal{E}}(X_{\mathcal{S}},X_{\mathcal{E}},C)]},$ (8)
with partition function
$\mathcal{Z}_{\mathcal{S}}^{*}(\lambda,C,\beta)=\int\mathrm{d}X_{\mathcal{S}}\,e^{-\beta\,\mathcal{H}_{\beta}^{*}(X_{\mathcal{S}},\lambda,C)}$ (9)
and marginal
$P^{\mathcal{S}}(X_{\mathcal{S}};\lambda,C,\beta)=\frac{e^{-\beta\,\mathcal{H}_{\beta}^{*}(X_{\mathcal{S}},\lambda,C)}}{\mathcal{Z}_{\mathcal{S}}^{*}(\lambda,C,\beta)}.$ (10)
Applied to the endpoints, this yields
$P^{\mathcal{S}}_{0}(X_{\mathcal{S}},\beta)$ $=P^{\mathcal{S}}(X_{\mathcal{S}};\lambda(0),C(0),\beta),$ (11)
$P^{\mathcal{S}}_{t_{\mathrm{eq}}}(X_{\mathcal{S}},\beta)$ $=P^{\mathcal{S}}(X_{\mathcal{S}};\lambda(t_{\mathrm{eq}}),C(t_{\mathrm{eq}}),\beta).$ (12)
The HMF free energy
$F^{*}_{\mathcal{S}}(\lambda,C,\beta)=-\beta^{-1}\ln\mathcal{Z}_{\mathcal{S}}^{*}(\lambda,C,\beta)$ (13)
defines the free energy difference
$\Delta F^{*}_{\mathcal{S}}(\beta)=F^{*}_{\mathcal{S}}(\lambda(t_{\mathrm{eq}}),C(t_{\mathrm{eq}}),\beta)-F^{*}_{\mathcal{S}}(\lambda(0),C(0),\beta).$ (14)
Theorem 1 (Endpoint equalities for HMF free energy).
For two equilibrium endpoints at the same inverse temperature $\beta$,
$e^{-\beta\Delta F^{*}_{\mathcal{S}}(\beta)}=\frac{\bigl\langle e^{-\beta\Delta\mathcal{H}^{*}_{\beta}(X_{\mathcal{S}})}\bigr\rangle_{\mathcal{S}}}{1+\chi^{2}\bigl(P^{\mathcal{S}}_{t_{\mathrm{eq}}}\parallel P^{\mathcal{S}}_{0}\bigr)},$ (15)
and
$e^{+\beta\Delta F^{*}_{\mathcal{S}}(\beta)}=\bigl\langle e^{+\beta\Delta\mathcal{H}^{*}_{\beta}(X_{\mathcal{S}})}\bigr\rangle_{\mathcal{S}}.$ (16)
Here
$\langle\bullet\rangle_{\mathcal{S}}=\int\mathrm{d}X_{\mathcal{S}}\,\bullet\,P^{\mathcal{S}}_{t_{\mathrm{eq}}}(X_{\mathcal{S}},\beta),$ (17)
the HMF shift is $\Delta\mathcal{H}^{*}_{\beta}(X_{\mathcal{S}})=\mathcal{H}^{*}_{\beta}(X_{\mathcal{S}},\lambda(t_{\mathrm{eq}}),C(t_{\mathrm{eq}}))-\mathcal{H}^{*}_{\beta}(X_{\mathcal{S}},\lambda(0),C(0))$, and the chi-squared divergence between endpoint marginals is
$\chi^{2}\bigl(P^{\mathcal{S}}_{t_{\mathrm{eq}}}\parallel P^{\mathcal{S}}_{0}\bigr)=-1+\int\mathrm{d}X_{\mathcal{S}}\,\frac{\bigl(P^{\mathcal{S}}_{t_{\mathrm{eq}}}(X_{\mathcal{S}},\beta)\bigr)^{2}}{P^{\mathcal{S}}_{0}(X_{\mathcal{S}},\beta)}.$ (18)
The result established here coincides with the endpoint equality derived in [rahbar2025thermodynamic], where the protocol is implicitly present. The trajectory map $\mathcal{T}_{t}$ pushes the initial distribution $P_{0}$ forward to a family $P_{t}$, and its system projection defines the time-dependent marginal $P_{t}^{\mathcal{S}}$. We postulate asymptotic equilibration,
$\lim_{t\rightarrow t_{\mathrm{eq}}}P_{t}^{\mathcal{S}}(X_{\mathcal{S}},\beta)=P^{\mathcal{S}}_{t_{\mathrm{eq}}}(X_{\mathcal{S}},\beta),$ (19)
which guarantees that $\Delta F^{*}_{\mathcal{S}}$ is well defined for any nonequilibrium protocol connecting the two equilibrium endpoints, without constraining the driven dynamics to be microscopically reversible. The explicit pushforward expressions and a detailed discussion of this boundary condition are given in [PRE]. Under asymptotic equilibration, the endpoint equalities admit exact trajectory forms. Introducing the average over the initial canonical ensemble,
$\langle\bullet\rangle_{X_{0}}=\int\mathrm{d}X_{0}\,\bullet\,P_{0}(X_{0},\beta),$ (20)
we can obtain the following theorem.
Theorem 2 (Trajectory counterpart of the endpoint equalities).
The HMF free energy difference satisfies
$e^{-\beta\Delta F^{*}_{\mathcal{S}}(\beta)}=\frac{\Big\langle e^{-\beta\Delta\mathcal{H}^{*}_{\beta}\bigl(\mathcal{T}_{t_{\mathrm{eq}}}^{\mathcal{S}}(X_{0})\bigr)}\Big\rangle_{X_{0}}}{1+\chi^{2}\bigl(P^{\mathcal{S}}_{t_{\mathrm{eq}}}\parallel P^{\mathcal{S}}_{0}\bigr)},$ (21)
and
$e^{+\beta\Delta F^{*}_{\mathcal{S}}(\beta)}=\Bigl\langle e^{+\beta\Delta\mathcal{H}^{*}_{\beta}\bigl(\mathcal{T}_{t_{\mathrm{eq}}}^{\mathcal{S}}(X_{0})\bigr)}\Bigr\rangle_{X_{0}}.$ (22)
The passage from Eqs. (15)-(16) to Eqs. (21)-(22) follows by inserting the trajectory representation of $P^{\mathcal{S}}_{t_{\mathrm{eq}}}$ and exchanging limits and integrals under the asymptotic equilibration condition (see [PRE]).
Frozen coupling regime. The scenario of frozen coupling deserves particular attention. In a broad class of nonequilibrium work protocols — both theoretical and experimental — the system remains in fixed coupling (the interaction strength $C$ between $\mathcal{S}$ and $\mathcal{E}$ is kept constant throughout the process) while the control parameter $\lambda(t)$ is driven. This is precisely the regime underlying the standard derivation of the JE [jarzynski1997equilibrium, jarzynski1997nonequilibrium, jarzynski2004nonequilibrium], although the JE itself is agnostic to whether the protocol acts through $\lambda$ or through $C$. Within this common, physically motivated freezing scenario, the formal structure of our derivation simplifies considerably. The complementary frozen-driving viewpoint, together with its direct connection to free energy perturbation (FEP), is developed in detail in the companion article [PRE].
Corollary 1 (Frozen coupling: trajectory forms reduce to bare-system expressions).
For a fixed coupling protocol $C(t)\equiv C(0)$, the HMF increment along a trajectory reduces exactly to the difference of bare-system Hamiltonians,
$\Delta\mathcal{H}^{*}_{\beta}\bigl(\mathcal{T}_{t_{\rm eq}}^{\mathcal{S}}(X_{0})\bigr)$
$=\mathcal{H}_{\beta}^{*}\bigl(\mathcal{T}_{t_{\rm eq}}^{\mathcal{S}}(X_{0}),\lambda(t_{\rm eq}),C(0)\bigr)-\mathcal{H}_{\beta}^{*}\bigl(\mathcal{T}_{t_{\rm eq}}^{\mathcal{S}}(X_{0}),\lambda(0),C(0)\bigr)$
$=\mathcal{H}_{\mathcal{S}}\bigl(\mathcal{T}_{t_{\rm eq}}^{\mathcal{S}}(X_{0}),\lambda(t_{\rm eq})\bigr)-\mathcal{H}_{\mathcal{S}}\bigl(\mathcal{T}_{t_{\rm eq}}^{\mathcal{S}}(X_{0}),\lambda(0)\bigr),$ (23)
Equation (1) follows directly from Eq. (Exact fluctuation relation for open systems beyond the Jarzynski equality), for constant $C$, the environmental contribution inside the logarithmic term cancels between the two endpoints. Substituting Eq. (1) into the trajectory equalities (21)-(22) yields
$e^{-\beta\Delta F^{*}_{\mathcal{S}}(\beta)}$
$=\frac{\Big\langle e^{-\beta\mathcal{H}_{\mathcal{S}}\bigl(\mathcal{T}_{t_{\mathrm{eq}}}^{\mathcal{S}}(X_{0}),\lambda(t_{\mathrm{eq}})\bigr)}e^{+\beta\mathcal{H}_{\mathcal{S}}\bigl(\mathcal{T}_{t_{\mathrm{eq}}}^{\mathcal{S}}(X_{0}),\lambda(0)\bigr)}\Big\rangle_{X_{0}}}{1+\chi^{2}\bigl(P^{\mathcal{S}}_{t_{\mathrm{eq}}}\parallel P^{\mathcal{S}}_{0}\bigr)},$ (24)
and the complementary identity,
$e^{+\beta\Delta F^{*}_{\mathcal{S}}(\beta)}$
$=\Big\langle e^{+\beta\mathcal{H}_{\mathcal{S}}\bigl(\mathcal{T}_{t_{\mathrm{eq}}}^{\mathcal{S}}(X_{0}),\lambda(t_{\mathrm{eq}})\bigr)}e^{-\beta\mathcal{H}_{\mathcal{S}}\bigl(\mathcal{T}_{t_{\mathrm{eq}}}^{\mathcal{S}}(X_{0}),\lambda(0)\bigr)}\Big\rangle_{X_{0}}.$ (25)
It is important to mention that as soon as $C(t)$ is non-frozen, the cancellation in (1) no longer occurs and the full HMF structure must be retained.
Corollary 2 (Heat-work-feedback decomposition for frozen coupling).
We now discuss the thermodynamic structure hidden in Eqs. (1) and  (1). In the frozen coupling regimes, energetic changes in $\mathcal{S}$ arise from (i) external manipulation of $\lambda(t)$ and (ii) exchange with the environment through the microscopic evolution of $X_{\mathcal{E}}$. To inspect these contributions, we start from Eq. (1) and apply a simple algebraic insertion-subtraction of the initial energy at fixed protocol, $\mathcal{H}_{\mathcal{S}}\bigl(X_{\mathcal{S}}(0|X_{0}),\lambda(0)\bigr)$, which yields the following representation. For $C(t)\equiv C(0)$,
$\mathcal{H}_{\mathcal{S}}\bigl(\mathcal{T}_{t_{\rm eq}}^{\mathcal{S}}(X_{0}),\lambda(t_{\rm eq})\bigr)-\mathcal{H}_{\mathcal{S}}\bigl(\mathcal{T}_{t_{\rm eq}}^{\mathcal{S}}(X_{0}),\lambda(0)\bigr)$
$=\underbrace{\Big[\mathcal{H}_{\mathcal{S}}\bigl(X_{\mathcal{S}}(t_{\mathrm{eq}}|X_{0}),\lambda(t_{\mathrm{eq}})\bigr)-\mathcal{H}_{\mathcal{S}}\bigl(X_{\mathcal{S}}(0|X_{0}),\lambda(0)\bigr)\Big]}_{I(t_{\mathrm{eq}}|X_{0})}$
$-\underbrace{\Big[\mathcal{H}_{\mathcal{S}}\bigl(X_{\mathcal{S}}(t_{\mathrm{eq}}|X_{0}),\lambda(0)\bigr)-\mathcal{H}_{\mathcal{S}}\bigl(X_{\mathcal{S}}(0|X_{0}),\lambda(0)\bigr)\Big]}_{II(t_{\mathrm{eq}}|X_{0})}.$ (26)
Term $I$. Along each realization, the system follows the trajectory $X_{\mathcal{S}}(t|X_{0})$ in its phase space. The total time variation of the bare–system Hamiltonian $\mathcal{H}_{\mathcal{S}}(X_{\mathcal{S}}(t|X_{0}),\lambda(t))$ is obtained by applying the chain rule to its explicit dependence on both $X_{\mathcal{S}}$ and the control parameter $\lambda(t)$,
$\frac{\mathrm{d}}{\mathrm{d}t}\,\mathcal{H}_{\mathcal{S}}\bigl(X_{\mathcal{S}}(t|X_{0}),\lambda(t)\bigr)=\underbrace{\bigl(\nabla_{X_{\mathcal{S}}}\mathcal{H}_{\mathcal{S}}\bigr)\cdot\dot{X}_{\mathcal{S}}}_{\dot{Q}_{\mathcal{S}}}+\underbrace{\frac{\partial\mathcal{H}_{\mathcal{S}}}{\partial\lambda}\,\dot{\lambda}}_{\dot{W}_{\mathcal{S}}}.$ (27)
The first term describes the energy exchange with the environment along the actual phase-space trajectory and defines the instantaneous heat rate, $\dot{Q}_{\mathcal{S}}$, whereas the second term accounts for the parametric energy input due to the protocol and defines the instantaneous work rate, $\dot{W}_{\mathcal{S}}$. Following the standard conventions of stochastic energetics [talkner2020colloquium], their time integrals give
$W_{\mathcal{S}}(t_{\mathrm{eq}}|X_{0})$ $=\int_{0}^{t_{\mathrm{eq}}}{\rm d}t\,\frac{\partial\mathcal{H}_{\mathcal{S}}}{\partial\lambda}\,\dot{\lambda}(t),$ (28)
$Q_{\mathcal{S}}(t_{\mathrm{eq}}|X_{0})$ $=\int_{0}^{t_{\mathrm{eq}}}{\rm d}t\,\big(\nabla_{X_{\mathcal{S}}}\mathcal{H}_{\mathcal{S}}\big)\cdot\dot{X}_{\mathcal{S}}(t|X_{0}),$ (29)
so that integrating Eq. (27) over $t\in[0,t_{\mathrm{eq}}]$ yields the total increment
$I(t_{\mathrm{eq}}|X_{0})=W_{\mathcal{S}}(t_{\mathrm{eq}}|X_{0})+Q_{\mathcal{S}}(t_{\mathrm{eq}}|X_{0}).$ (30)
Term $II$. For a fixed control $\lambda(0)$, $II(t_{\mathrm{eq}}|X_{0})$ can be written as
$II(t_{\mathrm{eq}}|X_{0})=\int_{0}^{t_{\mathrm{eq}}}\mathrm{d}t\;\Big(\nabla_{X_{\mathcal{S}}}\mathcal{H}_{\mathcal{S}}(X_{\mathcal{S}},\lambda(0))\Big)\cdot\dot{X}_{\mathcal{S}}.$ (31)
Here the gradient is evaluated with respect to $\mathcal{H}_{\mathcal{S}}$ frozen at $\lambda(0)$, whereas the $\dot{X}_{\mathcal{S}}$ belongs to the driven trajectory generated by $\lambda(t)$. Consequently, $II$ is a reference functional that projects the driven generalized velocity ($\dot{X}_{\mathcal{S}}$) onto the generalized force field ($\nabla_{X_{\mathcal{S}}}\mathcal{H}_{\mathcal{S}}$) of the initial stationary protocol. Because the reference and realized dynamics belong to distinct protocol layers, $II$ acts as a feedback-like correction [sagawa2010generalized] without a definite thermodynamic notion. Combining Eqs. (30) and (31) with Eq. (2), the total bare-system increment becomes
$W_{\mathcal{S}}(t_{\mathrm{eq}}|X_{0})+Q_{\mathcal{S}}(t_{\mathrm{eq}}|X_{0})-II(t_{\mathrm{eq}}|X_{0}).$ (32)
Substituting Eq. (32) into Eqs. (1)-(1) yields the heat-work-feedback representation of the endpoint equalities,
$e^{-\beta\Delta F_{\mathcal{S}}^{*}(\beta)}$ $=\frac{\big\langle e^{-\beta[\,W_{\mathcal{S}}+Q_{\mathcal{S}}-II\,]}\big\rangle_{X_{0}}}{1+\chi^{2}(P^{\mathcal{S}}_{t_{\mathrm{eq}}}\parallel P^{\mathcal{S}}_{0})},$ (33)
$e^{+\beta\Delta F_{\mathcal{S}}^{*}(\beta)}$ $=\big\langle e^{+\beta[\,W_{\mathcal{S}}+Q_{\mathcal{S}}-II\,]}\big\rangle_{X_{0}}.$ (34)
The decomposition isolates three pathwise contributions: (i) $W_{\mathcal{S}}$, mechanical work due to driving; (ii) $Q_{\mathcal{S}}$, total heat exchanged during the whole process; and (iii) $II$, a reference (projection) functional obtained by contracting the driven velocity ($\dot{X}_{\mathcal{S}}$) with the $\lambda(0)$ force field ($\nabla_{X_{\mathcal{S}}}\mathcal{H}_{\mathcal{S}}$). Only when $\lambda$ is held fixed does $II$ coincide with a thermodynamic heat.
GJE connection. The GJE restores $e^{-\beta\Delta F}$ under feedback by augmenting work with mutual information, providing an operational measure of information as a thermodynamic resource [sagawa2010generalized]. Our identities achieve restoration without invoking measurement-feedback a priori and without assuming LDB. The pathwise correction $Q_{\mathcal{S}}-II$ compensates dynamical asymmetry during driving, and the ensemble overlap $1+\chi^{2}(P^{\mathcal{S}}_{t_{\rm eq}}\parallel P^{\mathcal{S}}_{0})$ compensates endpoint mismatch. Formally, our equality is broader in scope, it holds without feedback and without LDB, while its structure is fully compatible with the GJE [sagawa2010generalized].
Figure 1: (A) System free energy difference $\Delta F_{\mathcal{S}}^{*}$ versus the final control value $\lambda(t_{\rm eq})$: circles (JE), squares (Eq. (1)), triangles (Eq. (1)), solid line (exact). Ramp/post-ramp protocol: $\lambda(t)$ increases linearly from $\lambda(0)$ to the target over $t_{\lambda}$, then is held fixed. (B) Representative run for $\lambda(0)=1\!\to\!\lambda(t_{\rm eq})=4$. Center: $\lambda(t)$. Top/bottom: ensemble trajectories of $x(t)$ and $p_{x}(t)$ (dashed line marks $t_{\lambda}$). Left/right: canonical marginals $P^{\mathcal{S}}_{0}(x,p_{x})$ and $P^{\mathcal{S}}_{t_{\rm eq}}(x,p_{x})$. During the ramp, dynamics are non-Liouvillian with damping $(\gamma_{\rm nl}^{\mathcal{S}},\gamma_{\rm nl}^{\mathcal{E}})$, post-ramp relaxation is governed by the underdamped Langevin equations, whose solution converges to the canonical distribution of the composite system, thereby satisfying the asymptotic equilibration postulate, Eq. (19).
JE as limiting case. The central identity underlying all our derivations is
$e^{-\beta\Delta F_{\mathcal{S}}^{*}}=\frac{Z^{*}(\lambda(t_{\mathrm{eq}}),C(t_{\mathrm{eq}}),\beta)}{Z^{*}(\lambda(0),C(0),\beta)}.$ (35)
By the definition of the HMF, $Z_{\mathcal{S}+\mathcal{E}}(\lambda,C,\beta)=Z^{*}(\lambda,C,\beta)\,Z_{\mathcal{E}}(\beta)$, so Eq. (35) is equivalent to
$\frac{Z^{*}(\lambda(t_{\mathrm{eq}}),C(t_{\mathrm{eq}}),\beta)}{Z^{*}(\lambda(0),C(0),\beta)}=\frac{Z_{\mathcal{S}+\mathcal{E}}(\lambda(t_{\mathrm{eq}}),C(t_{\mathrm{eq}}),\beta)}{Z_{\mathcal{S}+\mathcal{E}}(\lambda(0),C(0),\beta)}.$ (36)
This equality, which corresponds to [Eq. (23) in [jarzynski2004nonequilibrium]], contains no reference to the dynamics of the composite system. All relations derived in this work—independent of coupling strength or dynamical assumptions—reduce to this equilibrium partition function ratio. The standard JE can be viewed as the specific realization of this identity under Hamiltonian evolution of the composite system [jarzynski2004nonequilibrium]. For completeness, the detailed derivation linking Eq. (36) to the conventional JE is provided in the Appendix, where we recast Jarzynski’s original argument using the present notation.
Validation. To validate the framework, we examine how the proposed equalities hold beyond the regime where the JE applies. The model construction is inspired by the two-stage thought experiment introduced by Jarzynski in his reply [jarzynski2004nonequilibrium] to Cohen’s critique [cohen2004note]. In his argument, the control parameter is first driven from $\lambda(0)$ to $\lambda(t_{\lambda})$ (stage 1) and then held fixed at $\lambda=\lambda(t_{\lambda})$ for relaxation (stage 2). Two observers — one stopping at $t_{\lambda}$ (out of equilibrium) and the other at $t_{\mathrm{eq}}$ (after relaxation) — record the same work, since no work is performed once the driving protocol is fixed, and calculate the same $\Delta F_{\mathcal{S}}^{*}$, defined as the equilibrium free energy difference between the canonical states associated with $\lambda(0)$ and $\lambda(t_{\lambda})=\lambda(t_{\mathrm{eq}})$ via the HMF partition functions. As Jarzynski emphasized [jarzynski2004nonequilibrium], ”whether or not we choose to include a relaxation stage has no bearing on the validity” of the JE. While this statement is entirely correct, it implicitly depends on a key assumption about the underlying dynamics. To probe this assumption directly, we construct a validation model consisting of a non-Liouvillian ramp (stage 1) that breaks phase-space volume preservation, followed by an underdamped Langevin relaxation (stage 2) that enforces the required equilibrium endpoint. In this scenario, the JE no longer holds because phase-space contraction during the ramp breaks the Liouvillian structure of the composite dynamics, preventing the equilibrium partition function ratio in Eq. (36) from reproducing the JE. By contrast, the same partition function ratio leads directly to the endpoint relations in Thm. 1, without any dynamical assumptions, and the trajectory form in Thm. 2 follows once asymptotic equilibration is imposed as a boundary condition. For a detailed discussion of the validation model and protocol, see the Appendix.
Conclusions: In this Letter, we establish exact fluctuation relations for open systems. Using the HMF, the free energy difference is expressed via exponential moments of the HMF shift and an endpoint chi-squared divergence (Thm. 1). The trajectory counterparts of that relation are obtained under the assumption of asymptotic equilibration (Thm. 2). In the frozen coupling regime the relations reduce to bare-system expressions (Cor. 1) with a clear heat-work-feedback decomposition (Cor. 2), and they include the JE and its feedback extensions (GJE) as limiting cases. Validation on a composite model with a non-Liouvillian ramp followed by underdamped Langevin relaxation shows that our endpoint and trajectory relations reproduce the exact free energy differences, demonstrating accuracy and practical applicability beyond the JE regime. The generality of this framework invites direct application to complex nonequilibrium settings where traditional FTs are fragile or inapplicable, including active matter [johnsrud2025fluctuation, bebon2025thermodynamics], biological processes [woodside2014reconstructing], complex molecular and chemical environments [leighton2025flow], and open quantum systems [beyer2025operational]. In these regimes, our equalities provide a practical route to reconstruct free energy landscapes from realistic driving protocols without imposing idealized dynamical models or weak coupling assumptions. When applied to situations in which the JE is valid, our relations provide an exact re-expression of the JE that makes explicit the roles of endpoint overlap and system-environment interaction. In this sense, our relations can be interpreted as an enriched form of the JE, providing diagnostic tools that clarify how practical failures of the JE can arise, for example due to insufficient endpoint overlap [schmidt2025free], rare-event sampling limitations [kuang2020gamma], or unaccounted strong-coupling effects [talkner2020colloquium]. Together, this framework can potentially offer concrete guidance for designing and refining experimental and simulation protocols [casert2024learning], for controlling dissipation and optimizing driving in soft and active matter [johnsrud2025fluctuation, bebon2025thermodynamics, engel2023optimal], for improving nonequilibrium binding free energy calculations and molecular design workflows [kuang2020gamma, baumann2021challenges, serra2025nonequilibrium, serra2025path], and for advancing the operation and characterization of strongly coupled quantum devices and engines [cangemi2024quantum].
Acknowledgements.
We gratefully acknowledge financial support by the DFG under Germany’s Excellence Strategy EXC 2089/2-390776260 (e-conversion).
End Matter
Appendix: We collect detailed proofs of the theorems and corollaries stated in the main text. Each result is reproduced in full mathematical form to ensure completeness and clarity of the derivations.
Figure 2: (A) Schematic of the composite $\mathcal{S}+\mathcal{E}$ at a fixed bath temperature $T=\beta^{-1}/k_{B}$. The control $\lambda(t)$ (schematic purple spring) and the coupling $C(t)$ (schematic black spring) are driven independently and become constant at times $t_{\lambda}$ and $t_{C}$, respectively; the composite system then relaxes to equilibrium at $t_{\rm eq}$. The system free energy difference $\Delta F_{\mathcal{S}}^{*}(\beta)$ refers to the canonical endpoints $(\lambda(0),C(0))\!\to\!(\lambda(t_{\rm eq}),C(t_{\rm eq}))$ and is defined through the HMF partition functions. (B) An initial microstate $X_{0}\!\sim\!P_{0}(X,\beta)$ on $\Gamma(0)$ evolves under the kinematic map $X(t|X_{0})=\mathcal{T}_{t}(X_{0})$ through the ramp and the post-ramp relaxation, yielding the asymptotic density $P_{t_{\rm eq}}(X,\beta)$ on $\Gamma(t_{\rm eq})$. The frozen coupling case, $C(t)\equiv C(0)$, is the regime used in our validation.
Figure 3: Same plotting conventions as the Fig. 1. During the ramp we set $\gamma_{\rm nl}^{\mathcal{S}}=\gamma_{\rm nl}^{\mathcal{E}}=0$.
Table 1: Numerical inputs for the validation runs (frozen coupling; non-Liouvillian ramp, under damped Langevin relaxation). All quantities are in reduced, dimensionless units.
Quantity Value
Mass ($m$) $1.0$
System–environment coupling $C$ $1.0$
$\beta$ $1.0$
System–environment coupling $C$ $1.0$
Environment frequency $\omega$ $1.0$
Initial protocol $\lambda(0)$ $1.0$
Ramp duration $t_{\mathrm{ramp}}$ $10.0$
Relaxation duration $t_{\mathrm{eq}}-t_{\lambda}$ $20.0$
Ramp time step $\Delta t$ $2\times 10^{-3}$
Relaxation time step $\Delta t_{\mathrm{relax}}$ $2\times 10^{-2}$
Non-Liouvillian drag (system) $\gamma_{\mathrm{nl}}^{\mathcal{S}}$ $20.0$
Non-Liouvillian drag (environment) $\gamma_{\mathrm{nl}}^{\mathcal{E}}$ $5.0$
Langevin damping (relaxation) $\gamma$ $2.0$
Initial ensemble size $N_{\mathrm{traj}}$ $30000$
Relax integrator BAOAB [kieninger2022gromacs]
JE as limiting case. Starting from the Eq. (36), and assuming the composite system evolves under Hamilton’s equations generated by $\mathcal{H}_{\mathcal{S}+\mathcal{E}}$. The instantaneous energy change along a trajectory $X(t|X_{0})$ obeys
$\frac{\mathrm{d}}{\mathrm{d}t}\,\mathcal{H}_{\mathcal{S}+\mathcal{E}}(X(t|X_{0}),\lambda(t),C(t))$
$=\underbrace{\nabla_{X}\mathcal{H}_{\mathcal{S}+\mathcal{E}}\cdot\dot{X}}_{\text{advection}}+\frac{\partial\mathcal{H}_{\mathcal{S}}}{\partial\lambda}\,\dot{\lambda}+\frac{\partial\mathcal{V}_{\mathcal{S}\mathcal{E}}}{\partial C}\,\dot{C}.$ (37)
The advection term vanishes because the Hamiltonian flow is symplectic, $\nabla_{X}\mathcal{H}_{\mathcal{S}+\mathcal{E}}\cdot\dot{X}=\{\mathcal{H}_{\mathcal{S}+\mathcal{E}},\mathcal{H}_{\mathcal{S}+\mathcal{E}}\}=0$. Integrating over $t\in[0,t_{\mathrm{eq}}]$ gives [jarzynski2004nonequilibrium]
$\mathcal{H}_{\mathcal{S}+\mathcal{E}}(X(t|X_{0}),\lambda(t),C(t))-\mathcal{H}_{\mathcal{S}+\mathcal{E}}(X(0|X_{0}),\lambda(0),C(0))$
$=\int_{0}^{t_{\mathrm{eq}}}{\rm d}t\,\Big(\frac{\partial\mathcal{H}_{\mathcal{S}}}{\partial\lambda}\dot{\lambda}+\frac{\partial\mathcal{V}_{\mathcal{S}\mathcal{E}}}{\partial C}\dot{C}\Big)=:W_{\lambda}+W_{C}.$ (38)
Using the above equation and Liouville’s theorem, which guarantees a unit Jacobian for the change of variables ($X_{0}\mapsto X$), Jarzynski [jarzynski2004nonequilibrium] demonstrated that the equilibrium partition function ratio for the composite system can be expressed as
$\Big\langle e^{-\beta(W_{\lambda}+W_{C})}\Big\rangle_{X_{0},\beta}=\frac{Z_{\mathcal{S}+\mathcal{E}}(\lambda(t_{\mathrm{eq}}),C(t_{\mathrm{eq}}),\beta)}{Z_{\mathcal{S}+\mathcal{E}}(\lambda(0),C(0),\beta)}.$ (39)
For frozen coupling $C(t)\equiv C(0)$, Eq. (39) reduces to
$e^{-\beta\Delta F_{\mathcal{S}}^{*}}=\Big\langle e^{-\beta W_{\lambda}}\Big\rangle_{X_{0},\beta},$ (40)
which is the conventional JE expressed in the HMF notation. $\square$
Validation model and protocol: Model specification. We validate Eqs. (1) and (1) on an analytically tractable composite system consisting of a system with double well potential [whitelam2025improving, hanggi1990reaction, li2021equilibrium, saito1976relaxation, sun2003equilibrium] coupled to a harmonic environment. The system Hamiltonian is $\mathcal{H}_{\mathcal{S}}(x,p_{x};\lambda)=\frac{1}{2m}p_{x}^{2}+U_{\mathcal{S}}(x;\lambda)$, where $U_{\mathcal{S}}(x;\lambda)=\tfrac{1}{4}(x^{2}-\lambda)^{2}$, the environment Hamiltonian $\mathcal{H}_{\mathcal{E}}(y,p_{y})=\frac{1}{2m}p_{y}^{2}+\tfrac{1}{2}\omega^{2}y^{2}$, and the interaction $\mathcal{V}_{\mathcal{S}\mathcal{E}}(x,y;C)=Cxy$ with coupling constant $C$. The total Hamiltonian follows Eq. (Exact fluctuation relation for open systems beyond the Jarzynski equality). Eq. (Exact fluctuation relation for open systems beyond the Jarzynski equality) gives $\mathcal{H}_{\beta}^{*}(x,p_{x};\lambda,C)$, the corresponding marginal and partition function follow from $P^{\mathcal{S}}(X_{\mathcal{S}},\beta)={e^{-\beta\,\mathcal{H}_{\beta}^{*}(X_{\mathcal{S}},\lambda,C)}}/{\mathcal{Z}_{\mathcal{S}}^{*}(\lambda,C,\beta)}$ and $\mathcal{Z}_{\mathcal{S}}^{*}(\lambda,C,\beta)=\int\mathrm{d}X_{\mathcal{S}}\,e^{-\beta\,\mathcal{H}_{\beta}^{*}(X_{\mathcal{S}},\lambda,C)}$. The composite system is initialized in the canonical ensemble at $(\lambda(0),C(0))$, ensuring exact sampling from Eq. (11) (see Fig. 2). Two-stage protocol. We work in the frozen-coupling regime $C(t)\equiv C(0)$ and vary only $\lambda(t)$ (see Fig. 2). During the ramp $0\leq t\leq t_{\lambda}$, $\lambda(t)$ changes linearly from $\lambda(0)$ to $\lambda(t_{\lambda})$ (see Fig. 1). The composite evolves under non-Liouvillian deterministic dynamics $\dot{x}=p_{x}/m,\;\dot{p}_{x}=-\partial_{x}U_{\mathcal{S}}(x;\lambda(t))-Cy-\gamma_{\mathrm{nl}}^{\mathcal{S}}p_{x}/m,\;\dot{y}=p_{y}/m,\;\dot{p}_{y}=-\omega^{2}y-Cx-\gamma_{\mathrm{nl}}^{\mathcal{E}}p_{y}/m$, where $\gamma_{\mathrm{nl}}^{\mathcal{S}}$ and $\gamma_{\mathrm{nl}}^{\mathcal{E}}$ denote constant damping coefficients used only during the ramp. At $t=t_{\lambda}$, $\lambda(t)$ is held fixed and the composite relaxes under underdamped Langevin dynamics [hanggi1990reaction] obeying the FDT: $\dot{x}=p_{x}/m,\;\dot{p}_{x}=-\partial_{x}U_{\mathcal{S}}(x;\lambda(t_{\mathrm{eq}}))-Cy-\gamma p_{x}/m+\eta_{x}(t),\;\dot{y}=p_{y}/m,\;\dot{p}_{y}=-\omega^{2}y-Cx-\gamma p_{y}/m+\eta_{y}(t)$, where $\gamma$ is the friction coefficient [zwanzig2001nonequilibrium]. The terms $\eta_{x}(t)$ and $\eta_{y}(t)$ are independent standard Gaussian white noises with, $\langle\eta_{i}(t)\rangle=0$ and $\langle\eta_{i}(t)\eta_{j}(t^{\prime})\rangle=(2\gamma/\beta)\delta_{ij}\delta(t-t^{\prime})$ for $i,j\in\{x,y\}$. These equations converge to the canonical distribution associated with the frozen Hamiltonian $\mathcal{H}_{\mathcal{S}+\mathcal{E}}(x,p_{x},y,p_{y};\lambda(t_{\mathrm{eq}}),C)$, satisfying the asymptotic equilibration postulate Eq. (19). Estimators. After relaxation, the relations Eqs. (1) and (1) are evaluated. The ”Exact” curve is the equilibrium free energy difference from Eq. (35). For the JE estimator we accumulate the ramp work
$W_{\lambda}=\int_{0}^{t_{\lambda}}\partial_{\lambda}U_{\mathcal{S}}(x_{t};\lambda_{t})\,\dot{\lambda}\,\mathrm{d}t,$ (41)
with frozen $C$, and compute, $e^{-\beta\Delta F_{\mathcal{S}}^{*}}=\langle e^{-\beta W_{\lambda}}\rangle$. Fig. 1 reports $\Delta F_{\mathcal{S}}^{*}$ as a function of $\lambda(t_{\mathrm{eq}})$: the JE estimator deviates because the Liouvillian assumption is violated, whereas the endpoint equalities reproduce the exact equilibrium free energies. As a consistency check, when we set $\gamma_{\rm nl}^{\mathcal{S}}=\gamma_{\rm nl}^{\mathcal{E}}=0$ (Hamiltonian/Liouvillian ramp), the JE work estimator, the endpoint equalities, and the equilibrium reference all coincide, recovering the standard JE (see Fig.3). For completeness, all numerical parameters used for the simulations are listed in Table. 1.